# Communication-Efficient and Tensorized Federated Fine-Tuning of Large Language Models

### v1

@@ Abstract
Parameter-efficient fine-tuning (PEFT) methods typically assume that Large Language Models (LLMs) are trained on data from a single device or client. However, real-world scenarios often require fine-tuning these models on private data distributed across multiple devices. Federated Learning (FL) offers an appealing solution by preserving user privacy, as sensitive data remains on local devices during training. Nonetheless, integrating PEFT methods into FL introduces two main challenges: communication overhead and data heterogeneity. In this paper, we introduce FedTT and FedTT+, methods for adapting LLMs by integrating tensorized adapters into client-side models’ encoder/decoder blocks. FedTT is versatile and can be applied to both cross-silo FL and large-scale cross-device FL. FedTT+, an extension of FedTT tailored for cross-silo FL, enhances robustness against data heterogeneity by adaptively freezing portions of tensor factors, further reducing the number of trainable parameters. Experiments on BERT and LLaMA models demonstrate that our proposed methods successfully address data heterogeneity challenges and perform on par or even better than existing federated PEFT approaches while achieving up to 10$\times$ reduction in communication cost.
@@ 1 Introduction
Large Language Models (LLMs) like GPT-4 [1], LLaMA [2], and BERT [3] have become transformative tools in AI, excelling at tasks like translation and summarization [4], thanks to their advanced architectures such as Transformers [5]. These models can also be fine-tuned for specialized tasks using task-specific data [6], making them highly adaptable for various applications. However, due to the billions of parameters in modern LLMs, fully fine-tuning these models is computationally expensive and risks overfitting. To mitigate this, parameter-efficient fine-tuning (PEFT) methods like Adapters [7], Prompt-Tuning [8], and LoRA [9] have been developed, optimizing only a subset of parameters while freezing the majority, thus reducing computational costs without sacrificing performance [10].
While PEFT methods typically assume LLMs are trained on data from a single machine or client, real-world applications often require fine-tuning on private data distributed across multiple devices, such as sensitive medical records or legal documents [11, 12]. Privacy concerns make central data collection infeasible, highlighting the need for fine-tuning techniques that can adapt LLMs at the edge while preserving privacy. In this context, Federated Learning (FL) emerges as a compelling solution, as it prioritizes user data privacy by ensuring sensitive information remains on individual devices during training. Instead of transmitting data to a central server, clients in FL locally update their model parameters and share only model information, such as parameters or gradients, which are then aggregated by the server into a global model [13].
@@
To address the aforementioned problems, FL has been integrated into PEFT methods  [14, 15, 16]. These methods, however, often result in two major challenges: (i) communication and computation overhead in the FL system, and (ii) significant accuracy degradation, particularly under heterogeneous scenarios.
@@
Communication efficiency is crucial in FL, as edge devices typically have limited storage and computational power. While some methods attempt to reduce the number of training parameters by incorporating sparsity in PEFT approaches, they either perform poorly in FL [17, 18] or are computationally expensive for clients [19], rendering them unsuitable for practical scenarios.
@@
Data heterogeneity happens when training data is not identically and independently distributed across clients (non-i.i.d.). In such scenarios, local models on individual clients can diverge from the global model’s optimal state, leading to slower convergence [20, 21]. This issue is particularly pronounced when training LLMs in federated environments, as data heterogeneity can severely impact model performance [14]. Although various studies have proposed methods to address this challenge in federated learning [22, 23], these approaches often become inefficient in terms of communication and computational demands when applied to federated LLMs. Several works have modified LoRA to improve its efficiency in highly heterogeneous federated settings [24, 25]. However, these methods require communicating a large number of parameters (at least as much as LoRA).
@@
In this work, we present a federated and tensorized framework that reduces communication load compared to other PEFT methods, while maintaining comparable or even superior accuracy, and addressing challenges such as data heterogeneity. We introduce the Federated Tensor Train (FedTT) algorithm, where tensorized adapters serve as trainable parameters embedded within the encoder/decoder blocks of models used by clients. FedTT is applicable to both cross-silo FL, where all clients participate in training, and large-scale cross-device FL, where only a subset of clients is selected in each round. For cross-silo FL, we propose the FedTT+ algorithm, a heterogeneity-robust method that further reduces the number of parameters by adaptively freezing portions of the tensor factors in the models. We summarize our contributions as follows:
@@
  • We propose FedTT, an approach that effectively fine-tunes LLMs using tensorized adapters, achieving up to 10×10$\times$10 × less communication overhead compared to other popular federated PEFT methods.
  • We introduce FedTT+, an enhancement of FedTT that further reduces the number of trainable parameters and improves robustness against data heterogeneity in cross-silo FL. We demonstrate that FedTT+ outperforms state-of-the-art (SOTA) cross-silo FL methods with fewer trainable parameters.
  • We conduct extensive experiments with widely-used LLMs including the BERT and LLaMA-2 model families to validate the effectiveness of the proposed algorithms in reducing communication load while maintaining high accuracy.
@@ 2.1 Parameter Efficient Fine-Tuning (PEFT)
PEFT methods can be generally divided into three categories [26]. The first is Additive PEFT, where a small set of trainable parameters is added to the model, and only these are updated during training. Methods like Serial Adapter [7], Parallel Adapter [27], Prefix-tuning [28], and Prompt-tuning [8] fall into this category, and our approach aligns with this strategy. The second is Selective PEFT, which chooses a subset of existing model parameters for tuning, as seen in techniques like BitFit [29], U-Diff pruning [30], and PaFi [31]. Lastly, Reparameterized PEFT introduces a low-rank parameterization of pre-trained weights for training, with methods such as LoRA [9] and DoRA [32].
@@ 2.2 PEFT in Federated Setting
[14, 15] test and compare various PEFT methods such as Adapter, LoRA, Prompt Tuning, and BitFit in FL. Several works have modified LoRA to improve its efficiency in highly heterogeneous federated settings. For instance, SLoRA [24, 25] modify the initialization to address data heterogeneity, whereas HetLoRA [33] and FlexLoRA [34] adaptively adjust LoRA ranks for each client to handle system heterogeneity. In a recent work, FLoRA [35] slack LoRA matrices $A$ and $B$ for all clients and multiply two resulted matrices to prevent the interference resulting from FedAvg algorithm. However, these methods require communicating a large number of parameters (at least as much as LoRA).
Recently, [19] introduced sparse fine-tuning to reduce the communication load in federated LoRA. While this method lowers communication overhead and shows robustness to data heterogeneity in some tasks, it suffers from computational inefficiency due to the extensive matrix computations required during each communication round, both on the server and client sides. Additionally, their algorithm and results are limited to large-scale cross-device FL scenarios. FFA-LoRA [36] and RoLoRA [37] aim to improve accuracy in the presence of heterogeneity while simultaneously reducing the number of trainable parameters. Our algorithms are not only efficient in terms of communication overhead but also achieve comparable or better accuracy to these approaches particularly when data heterogeneity exists.
@@ 2.3 Tensor-based Model Compression
Recent advances in tensor compression have highlighted its potential to reduce model size and enhance training efficiency. Early work by novikov2015tensorizing [38] introduced the TT format, which represents weight matrices as sequences of tensor factors, laying the groundwork for subsequent developments. Recent efforts in tensorized fine-tuning of LLMs have shown promise in achieving high performance with significantly fewer parameters compared to traditional fine-tuning methods. However, these approaches still involve training a large number of variables, especially when compared to the more recent methods like LoRA. In response, yang2024loretta [39] developed the Low-Rank Economic Tensor-Train Adaptation (LoRETTA), which innovates by using tensorized-layer based adapters (LoRETTAadp) and reshaping update matrices into smaller tensor factors (LoRETTArep). Despite demonstrating substantial reductions in trainable parameters—up to 100 times less than popular PEFT methods—the performance of these tensorized approaches in FL scenarios, particularly in the presence of data heterogeneity, remains an open question.
@@ 3 Preliminaries
Figure 1: Illustration for the tensorized linear layer (a), FedTT algorithm (b), and the optional tensorized classifier applied for classification tasks (c). Note that we keep the model’s head uncompressed for language modeling tasks. The FedTT algorithm workflow includes fine-tuning clients’ tensorized adapters, aggregating tensor factors on the server, and broadcasting the updated weights back to clients.
@@ 3.1 Federated Fine-tuning
$\min_{\bm{w}}f(\tilde{\bm{w}},\bm{w})=\frac{1}{N}\sum_{i=1}^{N}f_{i}(\tilde{\bm{w}},\bm{w}),$
where $\tilde{\bm{w}}$ is the pre-trained model parameters, which are identical and fixed for all clients, and $f_{i}(\cdot)$ represents the local objective, expressed as:
$f_{i}(\tilde{\bm{w}},\bm{w})=\ell_{i}(\tilde{\bm{w}},\bm{w};\mathcal{D}_{i}),$
where $\ell_{i}(\cdot)$ is the local loss function, and $\mathcal{D}_{i}$ represents the local data distribution for client $c_{i}$.
A well-known FL method is the FedAvg algorithm [13]. The server selects a subset of clients $\mathcal{S}\subseteq\mathcal{C}$ in each training round. Each chosen client $c_{s}\in\mathcal{S}$ initializes its local model with the global model parameters from the previous round $\bm{w}^{(t)}$, and then performs local training using stochastic gradient descent on its dataset for $K$ local updates:
$\bm{w}_{s}^{(t)+k+1}\longleftarrow\bm{w}_{s}^{(t)+k}-\eta\nabla f_{s}(\tilde{\bm{w}},\bm{w}_{s}^{(t)+k}),$
where $\eta$ is the learning rate and $\bm{w}_{s}^{(t)+k}$ refers to the local model parameters for client $c_{s}$ during communication round $t$ and local update $k$. After local training is completed, clients send their updated model parameters to the server, which then aggregates these updates to form the new global model:
$\boldsymbol{w}^{(t+1)}=\frac{1}{|\mathcal{S}|}\sum_{s=1}^{|\mathcal{S}|}\boldsymbol{w}_{s}^{(t)+K}.$
Despite its advantages, FL faces two major challenges: (i) the large size of local models, which results in significant communication overhead, and (ii) data heterogeneity, which can cause local models to diverge from each other (client drift).
@@ 3.2 Tensor-Train Decomposition
In this subsection, we provide some introduction to the tensor and tensor train (TT) decomposition [40]. Tensors are natural multidimensional generalizations of matrices. The tensor $\mathcal{W}\in\mathbb{R}^{k_{1}\times\cdots\times k_{J}}$ is indexed as $\mathcal{W}=(w_{i_{1}\cdots i_{J}})_{1\leq i_{j}\leq k_{j}}$ said to have order of $J$ and dimension $k_{1},\cdots,k_{J}$. Given two tensors $\mathcal{W}\in\mathbb{R}^{k_{1}\times\cdots\times k_{J}}$ and $\mathcal{V}\in\mathbb{R}^{l_{1}\times\cdots\times l_{M}}$ with $k_{s}=l_{t}$, the multiplication between two tensors $\mathcal{C}=\mathcal{W}\times_{s,t}\mathcal{V}$ can be performed as:
@@
The TT decomposition serves as a potent alternative to traditional matrix decomposition techniques, which decompose a large tensor into a list of tensor factors [40] by TT-SVD method. As we show in Fig. 1 (a), to decompose the weight matrix into small tensors, it begins by reshaping a matrix $\boldsymbol{W}\in\mathbb{R}^{P\times Q}$ into a higher-dimensional tensor $\mathcal{W}\in\mathbb{R}^{k_{1}\times\cdots\times k_{J}}$. Then $\mathcal{W}$ can be parameterized compactly via a sequence of $J$ tensor factors $\mathcal{G}_{1},\ldots,\mathcal{G}_{J}$ as:
@@
Here, each tensor factor $\mathcal{G}_{j}$ has the shape of $\mathcal{G}_{j}\in\mathbb{R}^{r_{j-1}\times k_{j}\times r_{j}}$, where $\boldsymbol{r}=(r_{0},r_{1},\cdots,r_{J})$ is the tensor rank, and the product of dimensions $\Pi_{j}k_{j}=P\cdot Q$. The setup of tensor ranks follows the boundary conditions with $r_{0}=r_{J}=1$, while the other ranks $r_{j},j\notin{0,J}$ are chosen based on the specific tasks or made adaptive [41].
Instead of performing TT decomposition on the weight matrix, we directly initialize, store, and update the list of tensor factors in this work, as shown in Fig. 1 (a). During the forward pass, the tensor factors are directly contracted with the vector of activation values, and the weight matrix $\boldsymbol{W}$ does not need to be reconstructed. Since the size of the tensor factors is small, the contraction process is significantly faster than the original matrix-vector product.

### v2

@@ Abstract
Parameter-efficient fine-tuning (PEFT) methods typically assume that Large Language Models (LLMs) are trained on data from a single device or client. However, real-world scenarios often require fine-tuning these models on private data distributed across multiple devices. Federated Learning (FL) offers an appealing solution by preserving user privacy, as sensitive data remains on local devices during training. Nonetheless, integrating parameter-efficient fine-tuning (PEFT) methods into FL introduces two main challenges: communication overhead and data heterogeneity. In this paper, we introduce FedTT and FedTT+, methods for adapting LLMs by integrating tensorized adapters into client-side models’ encoder/decoder blocks. FedTT is versatile and can be applied to both cross-silo FL and large-scale cross-device FL. FedTT+, an extension of FedTT tailored for cross-silo FL, enhances robustness against data heterogeneity by adaptively freezing portions of tensor factors, further reducing the number of trainable parameters. Experiments on BERT and LLaMA family models demonstrate that our proposed methods successfully address data heterogeneity challenges and perform on par or even better than existing federated PEFT approaches while achieving up to 10$\times$ reduction in communication cost.
Communication-Efficient and Tensorized Federated Fine-Tuning of Large Language Models
Sajjad Ghiasvand1   Yifan Yang2   Zhiyu Xue1 Mahnoosh Alizadeh1   Zheng Zhang1   Ramtin Pedarsani1 Electrical and Computer Engineering Department, UC Santa Barbara1 Computer Science Department, UC Santa Barbara2 {sajjad,yifanyang,zhiyuxue,alizadeh,zhengzhang,ramtin}@ucsb.edu
@@ 1 Introduction
Large Language Models (LLMs) such as GPT-4 Achiam et al. (2023), LLaMA Touvron et al. (2023), and BERT Devlin et al. (2018) excel in tasks such as translation and summarization Bommasani et al. (2021) due to the capabilities of transformer architectures Vaswani (2017). While fine-tuning enhances adaptability Howard and Ruder (2018), fully fine-tuning these massive models is computationally expensive and prone to overfitting. Parameter-efficient fine-tuning (PEFT) methods such as Adapters Houlsby et al. (2019), Prompt-Tuning Lester et al. (2021), and LoRA Hu et al. (2021) address this by optimizing only a subset of parameters, reducing costs without sacrificing performance Ding et al. (2023). However, traditional PEFT assumes centralized data, whereas real-world applications often involve private, distributed data, such as medical or legal records Manoel et al. (2023); Shoham and Rappoport (2023); Soltanmohammadi and Hikmet (2024). Federated Learning (FL) emerges as a compelling solution, as it prioritizes user data privacy by ensuring sensitive information remains on individual devices during training. Instead of transmitting data to a central server, clients in FL locally update their model parameters and share only model information, such as parameters or gradients, which are then aggregated by the server into a global model McMahan et al. (2017a).
@@
To address the aforementioned problems, FL has been integrated into PEFT methods  Zhang et al. (2023); Fan et al. (2023); Zhao et al. (2023). These methods, however, often result in two major challenges: (i) communication and computation overhead in the FL system, and (ii) significant accuracy degradation, particularly under heterogeneous scenarios.
@@
Communication efficiency is crucial in FL, as edge devices typically have limited storage and computational power. While some methods attempt to reduce the number of training parameters by incorporating sparsity in PEFT approaches, they either perform poorly in FL He et al. (2022); Wu and Chen (2022) or are computationally expensive for clients Kuo et al. (2024), rendering them unsuitable for practical scenarios.
@@
Data heterogeneity happens when training data is not identically and independently distributed across clients (non-i.i.d.). In such scenarios, local models on individual clients can diverge from the global model’s optimal state, leading to slower convergence Hsieh et al. (2020); Li et al. (2020). This issue is particularly pronounced when training LLMs in federated environments, as data heterogeneity can severely impact model performance Zhang et al. (2023). Although various studies have employed techniques such as gradient tracking to address this challenge in federated learning Ghiasvand et al. (2024); Ebrahimi et al. (2024), these approaches often become inefficient in terms of communication and computational demands when applied to federated LLMs. Several works have modified LoRA to improve its efficiency in highly heterogeneous federated settings Babakniya et al. (2023). However, these methods require communicating a large number of parameters (at least as much as LoRA).
@@
In this work, we present a federated and tensorized framework that reduces communication load compared to other PEFT methods, while maintaining comparable or even superior accuracy, and addressing challenges such as data heterogeneity. We introduce the Federated Tensor Train (FedTT) algorithm, where tensorized adapters serve as trainable parameters embedded within the encoder/decoder blocks of models used by clients. FedTT is applicable to both cross-silo FL, where all clients participate in training, and large-scale cross-device (LSCD) FL, where only a subset of clients is selected in each round. For cross-silo FL, we propose the FedTT+ algorithm, a heterogeneity-robust method that further reduces the number of parameters by adaptively freezing portions of the tensor factors in the models. We summarize our contributions as follows:
@@
  • We propose FedTT, an efficient FL method to fine-tune LLMs using tensorized adapters, achieving up to 10×10$\times$10 × communication reduction than other popular federated PEFT methods.
  • We introduce FedTT+, an enhancement of FedTT that further reduces the number of trainable parameters and improves robustness against data heterogeneity in cross-silo FL. FedTT+ outperforms state-of-the-art (SOTA) cross-silo FL methods with fewer trainable parameters.
  • We conduct extensive experiments across various settings, including data heterogeneity and differential privacy, using widely adopted LLMs such as the BERT and LLaMA-2 model families. These experiments validate the effectiveness of our proposed algorithms in reducing communication overhead while preserving high accuracy.
@@ 2.1 Parameter Efficient Fine-Tuning (PEFT)
PEFT methods can be generally divided into three categories Han et al. (2024). The first is Additive PEFT, where a small set of trainable parameters is added to the model, and only these are updated during training. Methods like Prefix-tuning Li and Liang (2021) and Prompt-tuning Lester et al. (2021) fall into this category, and our approach aligns with this strategy. The second is Selective PEFT, which chooses a subset of existing model parameters for tuning, as seen in techniques like BitFit Zaken et al. (2021) and PaFi Liao et al. (2023). Lastly, Reparameterized PEFT introduces a low-rank parameterization of pre-trained weights for training, with methods such as LoRA Hu et al. (2021) and DoRA Liu et al. (2024).
@@ 2.2 PEFT in Federated Setting
(Zhang et al., 2023) tests and compares various PEFT methods such as Adapter, LoRA, Prompt Tuning, and BitFit in FL. Several works have modified LoRA to improve its efficiency in highly heterogeneous federated settings. For example, SLoRA Babakniya et al. (2023); Yan et al. (2024) modify initialization to address data heterogeneity, whereas HetLoRA Cho et al. (2023) and FlexLoRA Bai et al. (2024) adaptively adjust LoRA ranks for each client to handle system heterogeneity. However, these methods still require communicating massive parameters.
Recently, Kuo et al. (2024) introduced sparse fine-tuning to reduce the communication load in federated LoRA. Although this method lowers communication overhead and shows robustness to data heterogeneity in some tasks, it suffers from computational inefficiency due to the extensive matrix computations required during each communication round, both on the server and client sides. FFA-LoRA Sun et al. (2024) and RoLoRA Chen et al. aim to improve accuracy in the presence of heterogeneity while simultaneously reducing trainable parameters. Our algorithms can reduce communication overhead while achieving comparable or better accuracy to these approaches, particularly when data heterogeneity exists.
Kim et al. (2023) employs hypernetworks for adapters to reduce the number of trainable parameters, and this approach can be integrated with any PEFT method, including ours. However, it introduces additional computational overhead, as each client must generate adapters from the hypernetwork Hu et al. (2024).
@@ 2.3 Tensor-based Model Compression
Tensor compression has shown its great potential to reduce model size and enhance training efficiency. Early work by Novikov et al. (2015) employeed the TT format for network compression. Recent efforts in tensorized fine-tuning of LLMs have shown promise in achieving high performance with significantly fewer parameters compared to traditional fine-tuning methods. However, these approaches still involve training a large number of variables, especially when compared to the more recent methods like LoRA. In response, Yang et al. (2024b) developed the Low-Rank Economic Tensor-Train Adaptation, which innovates by using tensorized-layer based adapters and reshaping update matrices into smaller tensor factors. Despite demonstrating substantial reductions in trainable parameters—up to 100 times less than popular PEFT methods—the performance of these tensorized approaches in FL scenarios, particularly in the presence of data heterogeneity, remains an open question.
@@ 3 Preliminaries
Figure 1: Illustration for the tensorized linear layer (a), FedTT algorithm (b), and the optional tensorized classifier applied for classification tasks (c). The FedTT algorithm workflow includes fine-tuning clients’ tensorized adapters, aggregating tensor factors on the server, and broadcasting the updated weights back to clients.
@@ 3.1 Federated Fine-tuning
$\min_{\bm{w}}\mathcal{L}(\tilde{\bm{w}},\bm{w})=\frac{1}{N}\sum_{i=1}^{N}\ell_{i}(\tilde{\bm{w}},\bm{w};\mathcal{D}_{i}),$
where $\tilde{\bm{w}}$ is the pre-trained model parameters, which are identical and fixed for all clients, $\ell_{i}(\cdot)$ represents the local objective, and $\mathcal{D}_{i}$ represents the local data distribution for client $c_{i}$.
A well-known FL method is the FedAvg algorithm McMahan et al. (2017a). The server selects a subset of clients $\mathcal{S}\subseteq\mathcal{C}$ in each training round. Each chosen client $c_{s}\in\mathcal{S}$ initializes its local model with the global model parameters from the previous round $\bm{w}^{(t)}$, and then performs local training using stochastic gradient descent on its dataset for $K$ local updates:
$\bm{w}_{s}^{(t)+k+1}\longleftarrow\bm{w}_{s}^{(t)+k}-\eta\nabla\ell_{s}(\tilde{\bm{w}},\bm{w}_{s}^{(t)+k};\mathcal{D}_{s}),$
where $\eta$ is the learning rate and $\bm{w}_{s}^{(t)+k}$ refers to the local model parameters for client $c_{s}$ during communication round $t$ and local update $k$. After local training is completed, clients send their updated model parameters to the server, which then aggregates these updates to form the new global model: $\bm{w}^{(t+1)}=1/|\mathcal{S}|\sum_{s=1}^{|\mathcal{S}|}\bm{w}_{s}^{(t)+K}.$
Despite its advantages, FL faces two major challenges: (i) the large size of local models, which results in significant communication overhead, and (ii) data heterogeneity, which can cause local models to diverge from each other.
@@ 3.2 Tensor-Train Decomposition
In this subsection, we provide some introduction to the tensor and tensor train (TT) decomposition Oseledets (2011). Tensors are natural multidimensional generalizations of matrices. The tensor $\mathcal{W}\in\mathbb{R}^{k_{1}\times\cdots\times k_{J}}$ is indexed as $\mathcal{W}=(w_{i_{1}\cdots i_{J}})_{1\leq i_{j}\leq k_{j}}$ said to have order of $J$ and dimension $k_{1},\cdots,k_{J}$. Given two tensors $\mathcal{W}\in\mathbb{R}^{k_{1}\times\cdots\times k_{J}}$ and $\mathcal{V}\in\mathbb{R}^{l_{1}\times\cdots\times l_{M}}$ with $k_{s}=l_{t}$, the multiplication between two tensors $\mathcal{C}=\mathcal{W}\times_{s,t}\mathcal{V}$ can be performed as:
@@
The TT decomposition serves as a potent alternative to traditional matrix decomposition techniques, which decompose a large tensor into a list of tensor factors Oseledets (2011) by TT-SVD method. As shown in Fig. 1 (a), to decompose the weight matrix into small tensors, we begin by reshaping a matrix $\bm{W}\in\mathbb{R}^{P\times Q}$ into a tensor $\mathcal{W}\in\mathbb{R}^{k_{1}\times\cdots\times k_{J}}$. Then $\mathcal{W}$ can be parameterized compactly via a sequence of $J$ tensor factors $\mathcal{G}_{1},\ldots,\mathcal{G}_{J}$ as:
@@
Here, each tensor factor $\mathcal{G}_{j}$ has the shape of $\mathcal{G}_{j}\in\mathbb{R}^{r_{j-1}\times k_{j}\times r_{j}}$, where $\bm{r}=(r_{0},r_{1},\cdots,r_{J})$ is the tensor rank, and the product of dimensions $\Pi_{j}k_{j}=P\cdot Q$. The setup of tensor ranks follows the boundary conditions with $r_{0}=r_{J}=1$, while the other ranks $r_{j},j\notin{0,J}$ are chosen based on specific tasks or made adaptive Yang et al. (2024c).
As we can see, the tensorized layer substantially reduces the parameter count for the weight matrix $\bm{W}$ from $P\times Q$ to $\sum_{i=1}^{d}r_{i-1}k_{i}r_{i}$, offering a much higher compression ratio than previous PEFT work. Unlike the traditional Adapters method Houlsby et al. (2019), which uses a bottleneck structure to reduce trainable parameters, our tensorized adapters achieve an even larger compression ratio by using two tensorized linear layers with a nonlinear activation in between. For example, with weight $\bm{W}$ with size $768\times 768$ and bottleneck size $64$, a standard Adapter incurs $2\cdot 768\cdot 64\approx 98K$ trainable parameters for its weight matrices, whereas our method adds only $\sum_{i=1}^{6}(5^{2}\cdot 8)=1.2K$ parameters (assuming core dimensions $[8,8,8,8,8,8]$ for $\mathcal{W}$ and TT rank 5). This high compression ratio enables FedTT to achieve higher performance under a similar communication cost compared with other PEFT approaches.
Instead of performing TT decomposition on the weight matrix, we directly initialize, store, and update the list of tensor factors in this work, as shown in Fig. 1 (a). During the forward pass, the tensor factors are directly contracted with the vector of activation values, and the weight matrix $\bm{W}$ does not need to be reconstructed. Since the size of the tensor factors is small, the contraction process is significantly faster than the original matrix-vector product Yang et al. (2024c).